Convert each of the following indirect quotes to dollar direct quotes: a. $1 = 20,864 Vietnamese dong (Round your answer to 7c. $1 = 9.175 South African rand (Round your answer to 4 decimal places.) Exchange rateſ
Answer #1

Solution :

a. As per the information given in the question

$1 = 20,864 Vietnamese dong

Thus the dollar direct quote of is

= 1 / 20,864

= 0.00004792944

= 0.0000479 ( when rounded off to seven decimal places )

Thus 1 Vietnamese dong = $ 0.0000479

b. As per the information given in the question

$1 = 6.300 Venezuelan bolivar fuerte

Thus the dollar direct quote is

= 1 / 6.300

= 0.158730159

= 0.1587 ( when rounded off to four decimal places )

Thus 1 Venezuelan bolivar fuerte = $ 0.1587

c. As per the information given in the question

$1 = 9.175 South African Rand

Thus the dollar direct quote is

= 1 / 9.175

= 0.108991826

= 0.1090 ( when rounded off to four decimal places )

Thus 1 South African Rand = $ 0.1090